SPECTROSCOPIC CHARATERIZATION OF ThF AND THE LOW-LYING STATES OF ThF^{+}

\underline{BEAU J. BARKER}, IVAN O. ANTONOV, and MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Theoretical calculations predict that internal electric fields as high as 90 GV/cm can be attained by polarizing the ^{3}Δ_{1} state of ThF^{+}^{a} Consequently, this ion is of interest for investigation of the dipole moment of the electron. However, spectroscopic data have not been reported previously for ThF or ThF^{+}. We have used laser induced fluorescence and resonantly enhanced two-photon ionization to examine ThF. Multiple electronic transitions were observed in the 19530-21300 cm^{−1}range. Rotationally resolved data have been obtained, and the ground state is shown to be X^{2}Δ_{3}_{/}_{2}. Pulsed field ionization - zero electron kinetic energy spectra have been recorded for the ThF+ cation. Vibronic progressions of the X^{1}Σ^{+} and excited ^{3}Δ states have been identified. The term energy for the ^{3}Δ_{1} state was found to be T_{0}=330 cm^{−1}. Details of the experiments and spectroscopic constants for ThF and ThF^{+} will be reported.

aE. R. Meyer and J. L. Bohn, Phys. Rev. A At., Mol., Opt. Phys. \underline{78}, 010502/1 (2008). ”Prospects for an electron electric-dipole moment search in metastable ThO and ThF+”